# Stochastic Re-weighted Gradient Descent via Distributionally Robust Optimization

### v4

@@ F.9 Convergence of RGD and additional costs
In the extreme class imbalance setting, we note that uniform sampling for mini-batch generation + RGD re-weighting (as done in Algorithm 1) would be slower to converge than using importance sampling for mini-batch generation. This is because the former tends to have higher variance. But this is easily fixable in Algorithm 1. We simply update the mini-batch generation step with importance sampling; that is, we select point ‘$i$’ with probability proportional to its weight $exp(l_{i})$ (instead of uniform sampling that is currently done). The main reason for not considering this in this work is our desire to illustrate the generality of our approach and its applicability to a wide variety of learning tasks, without focusing too much on the class imbalance task. We believe this generality and simplicity is what makes our method quite attractive to the practitioner as showcased in some of experiments including Natural Language Processing, Image Classification, Tabular Classification, Distribution Shifts, and Meta-learning. Furthermore, Figure 6 illustrates the convergence plots of miniGPT pre-training and ViT-S on Imagenet-1K. Overall, we note similar stable training convergence on both, while RGD is able to focus more heavily on harder samples and reach a better minima.

### v5

@@ D.9 Convergence of RGD and additional costs
In the extreme class imbalance setting, we note that uniform sampling for mini-batch generation + RGD re-weighting (as done in Algorithm 1) would be slower to converge than using importance sampling for mini-batch generation. This is because the former tends to have higher variance. But this is easily fixable in Algorithm 1. We simply update the mini-batch generation step with importance sampling; that is, we select point ‘$i$’ with probability proportional to its weight $\exp(\ell_{i})$ (instead of uniform sampling that is currently done). The main reason for not considering this in this work is our desire to illustrate the generality of our approach and its applicability to a wide variety of learning tasks, without focusing too much on the class imbalance task. We believe this generality and simplicity is what makes our method quite attractive to the practitioner as showcased in some of experiments including Natural Language Processing, Image Classification, Tabular Classification, Distribution Shifts, and Meta-learning. Furthermore, Figure 6 illustrates the convergence plots of miniGPT pre-training and ViT-S on Imagenet-1K. Overall, we note similar stable training convergence on both, while RGD is able to focus more heavily on harder samples and reach a better minima.